Consider the window (of area 0.34 m2) of an airplane. If the pressure inside the plane is atmospheric pressure, while the pressure outside is 24 % of Patm, find the magnitude of the force on the window. Ignore the velocity of the airplane (i.e., ignore Bernoulli’s effect).
Answer #1

Pahn - ২42 ab ?am १oc) = २५ १atm - 2५ x?arm Patom | 6০ 16 १atm 100 १atm= |o138Sa 76x101325 100 \o132ऽ foet= 77007 Pa %3D fnet
Answer #2

SOLUTION :


Inside pressure = 1 atmosphere = 101325 Pa


Outside pressure = 0.24 atmosphere = 101325 * 0.24 = 24318 Pa 


These pressures act in opposite directions.


So, net pressure on the window = 101325 - 24318 = 77007 Pa = 77007 N/m^2


So, 


force on the window 

= Net pressure  in N/m^2  * Area in m^2

= 770007 * 0.34  N

= 26182.38 N (ANSWER).

answered by: Tulsiram Garg